# Generating Enhanced Negatives for Training Language-Based Object Detectors

### v1

@@ Abstract
The recent progress in language-based open-vocabulary object detection can be largely attributed to finding better ways of leveraging large-scale data with free-form text annotations. Training such models with a discriminative objective function has proven successful, but requires good positive and negative samples. However, the free-form nature and the open vocabulary of object descriptions make the space of negatives extremely large. Prior works randomly sample negatives or use rule-based techniques to build them. In contrast, we propose to leverage the vast knowledge built into modern generative models to automatically build negatives that are more relevant to the original data. Specifically, we use large-language-models to generate negative text descriptions, and text-to-image diffusion models to also generate corresponding negative images. Our experimental analysis confirms the relevance of the generated negative data, and its use in language-based detectors improves performance on two complex benchmarks.
@@ 1 Introduction
Using natural language in object detection to describe semantics bears the potential to significantly increase the size of the detector’s label space and enable novel applications. While standard detectors operate on a fixed label space [37, 23, 41], natural language allows for a broad spectrum of object descriptions, ranging from generic terms like ”vehicle” to specific expressions like ”the red sports car parked on the left side” [40, 52, 12, 25, 53, 30, 17]. Several works advanced language-based object detection over the past few years with novel training strategies [6, 56, 4, 19, 22, 33] and model architectures [15, 32, 44, 11].
@@
Our experiments demonstrate clear accuracy gains on two challenging benchmarks, +2.9AP on OmniLabel [40] and +3.3AP on D [52], when adding our automatically-generated negative data into the training of baseline models like GLIP or FIBER. Moreover, we provide an in-depth analysis of the generated data (text and images) and how they contribute to better language-based detection.
@@ 2 Related work
Negative samples in object detection:
The notion of negatives is crucial for training discriminative models [38, 24]. Also for object detection, hard negative mining [43] has proven beneficial for model training. However, these prior works aim to find hard negative training examples rather than negatives in the label space, because the label space is fixed in standard detection. For language-based datasets, the space of potential negatives is extremely large because object descriptions are free-form text. Prior works [5, 29, 54, 45, 42] investigate negative texts for general vision & language models with different strategies, including changing individual words (foil) with rules based on knowledge graphs [31] or with LLMs. For language-based detection specifically, [8] explores such rule-based foils, while [20] uses LLMs with specific templates to replace object names with alternative descriptions. In contrast, our work (1) focuses on the localization task, (2) explores more comprehensive strategies to generate negatives with LLMs, and (3) proposes to also generate corresponding negative images with text-to-image diffusion models.
@@ 3.2.1 Generating negative descriptions with LLMs
Prior work [42, 5, 8] explored rule-based approaches to generate negative text. However, such rules are typically limited to simple knowledge graphs and are limited to replacing only individual words, often just nouns, or swapping words. In contrast, we explore more powerful LLMs to automatically generate relevant negatives. To make the negative text generation efficient, in all cases, we first leverage a strong instruction-tuned LLM [34] to generate 50k positive-negative pairs, and then finetune a LLaMA-7B [46] model with those pairs to then generate negative captions on large grounding datasets. In the following, we describe three ways to instruct an LLM for generating positive-negative pairs of object descriptions:
@@
In this approach, we first prompt an instruction-tuned LLM [34] to find concepts (i.e., objects, attributes and relationships) in object descriptions. Compared to rule-based parsers [50], LLMs can provide richer information. For example, for the caption “A transportation vehicle is carrying a crowd of people who are sitting and standing.”, the parser ignores “sitting” and “standing”, while LLMs regard them as attributes. Then, we pick one concept from the first step sequentially and prompt LLMs again to generate a negative caption by changing the concept. For both steps, the prompts are manually curated with the task definition and step-by-step instructions for the generation. Please find the exact prompt for the LLM in the supplement.
@@
Next, we attempt to give the LLM more freedom in generating negative descriptions. We first prompt it to identify all objects in the original caption, and then to re-combine them to create a new sentence different from the original one. We allow the LLM to ignore, change or add new objects. For example, given the caption “A boy is playing with his dog” and two objects “boy” and “dog”, the LLM can output “The girl and her dog are playing fetch in the park”. Detailed prompts for both identifying objects and re-combination are in the supplement.
@@
Finally, we attempt to extract text descriptions about the difference between positive and corresponding negative descriptions by using human-annotated positive-negative pairs. We randomly sample 80 pairs of positive and negative text from the Winoground dataset [45] and prompt the instruction-tuned LLM [34] to summarize the difference of those text pairs as texts. Then, instead of manually creating prompts to generate positive-negative pairs, we leverage the text summary together with three randomly sampled Winoground pairs as prompt to generate such data. This pipeline does not require hand-crafted prompts to explain the concept of negatives and how to create them to the LLM. LABEL:lst:prompt_incontextsummarygen shows abbreviated prompts for using the summary to generate positive-negative pairs. The supplement contains the full prompts for generating a summary and generating positive-negative pairs.
@@ 3.3 Learning from negative samples
We now describe how to integrate the generated negative text and images into training language-based detectors.
@@
The generated data does not prescribe any specific architecture for the detector. A common choice, which we also use for our experiments, is [19, 4]. The inputs are image and text , and the output is a set of bounding boxes with corresponding logits . Here, is the number of tokens required to represent text . The ground truth can be represented by a binary assignment matrix . Rows refer to ground truth boxes and columns to tokens in . Each element indicates if a token corresponds to a box , which is given by the ground truth indices . To define a loss, bipartite graph matching associates predictions with ground truth. For matched predictions, the target vector is the corresponding row from , while it is an all-zero vector for unmatched targets. The loss is then computed as . Fig. 4 illustrates the loss.
@@ 4.1 Experimental design
We choose two recently proposed benchmarks, OmniLabel [40] and D [52], as our test beds. These benchmarks evaluate more aspects of language-based detection than existing referring expressions [53, 30, 49] or open-vocabulary detection [17, 7] benchmarks. Specifically, both benchmarks contain complex object descriptions that go beyond simple category names from open-vocabulary detection benchmarks. Moreover, the descriptions can refer to zero, one or multiple instances in the image, in contrast to standard referring expression benchmarks. These properties enable a more stringent evaluation metric as in object detection, which is based on average precision (AP) in both OmniLabel [40] and D [52]. Both benchmarks provide more fine-grained metrics. OmniLabel evaluates separately for categories, descriptions, and descriptions referring to at least one object, with APc, APd and APd-P, respectively. D differentiates descriptions on absence (“Abs”) and presence (“Pres”) that indicate whether or not they contain any form of negation (e.g., “without”), as well as on text lengths. Finally, we create a specific split for OmniLabel, “OmniLabel-Negative”, to evaluate the model only on images that contain at least one negative description (i.e., not referring to any object).
@@
| D (default) | D (by length of texts) | ||||||
@@ 4.2 Benchmark comparisons
Tabs. 1 and 2 evaluate the impact of our generated negative training data on the OmniLabel [40] and D [52] benchmarks. In both tables, the first set of rows are baselines provided by the benchmarks. The following rows show the main comparisons for GLIP-T [19] and FIBER-B [4] with and without adding our generated negative training data. First, we can see that adding negative data improves results across all metrics for both models and both benchmarks. On OmniLabel, we can see a +2.9% and +2.4% increase in AP for GLIP-T and FIBER-B, respectively. Similarly, we observe a +2.3% and +3.3% increase in the main metric of D (AP on full descriptions) for GLIP-T and FIBER-B.
| Whole OmniLabel | OmniLabel-Negative | D | |||||||||
@@ 4.3 Analysis on negative texts
Effectiveness of different negative texts: We finetune FIBER-B without and with different kinds of negative texts mentioned in Sect. 3.2.1, i.e., Rule-based foils, LLM-based foils, Recombination with LLMs, In-context summary with LLMs, and present results in Table 4. We find all kinds of negatives improve the original FIBER-B on both OmniLabel and D benchmarks. Negative texts from LLMs generally achieve better results compared to LLM-based foils, which indicates that LLMs are powerful tools for negative text generation. Moreover, both recombination and in-context summary with LLMs outperform LLM-based foils in all metrics except APd-P. Note that APd-P refers to evaluations without negative label spaces, which is a task weaker than language-based detection. Based on such results, we argue that although word foils provide promising results in traditional studies [8, 42], it is sub-optimal to LLMs. We need to explore varied ways to unlock the ability of LLMs. We believe that our two solutions, i.e., recombination and in-context summary, provide a good starting point for future studies.
@@ 4.4 Analysis on negative images
Following the idea of concatenating the positive and negative captions as the text input, we concatenate the positive and negative images as one input image during training. See supplement for an example. In this way, models are forced to tell the difference between the positive and negative images within one training iteration, which helps detectors to learn better about the negative. As shown in Table 4, such a simple technique improves upon “+ Box&CLIP filters” both on OmniLabel and D. Furthermore, we ensemble the weights of two FIBER models, one finetuned with negative texts only, and the other finetuned with both negative texts and images. Finally, compared to using negative texts only, we gain 1.3 APd on OmniLabel and no performance drop on D.
@@
Table 4 shows that negative images help on OmniLabel but not much on D. We explore this on a data basis. We first crop image regions for generated images, OmniLabel images, and D images based on the bounding boxes. Then, we randomly select 1000 image regions and feed them into a CLIP image encoder to get CLIP embeddings. Later, we input those embeddings to t-SNE [47] to illustrate the similarities between different image regions. As shown in Fig. 11, D’s regions are grouped into several clusters, while OmniLabel and our generated regions are scattered in the center. This indicates that there is a clear domain gap between D and the other two. Thus, it is plausible that our generated images only helps on OmniLabel. In our view, the gap comes from that D collect data in groups based on categories. In contrast, OmniLabel collects data randomly.
@@ 6.3.1 Combining negative texts of different sources
In this part, we explore to combine negative texts of different sources. When two or more sources are combined, we randomly sample negative texts from the combined source during training. Table 5 provides the results of different combinations, from which we have the following findings. First, all combinations outperform the original FIBER in terms of almost all the metrics, which indicates the effectiveness of adding negative texts. Second, there is no combination that consistently outperforms others in all metrics. For example, “+ (4) Texts for negative images” achieves the leading performance on the whole OmniLabel but not on OmniLabel-Negative and D. Moreover, the performance gaps between different combinations are minor. Some possible explanations to such results are: 1) All four methods generate negative texts mainly by changing concepts of the original captions. Thus, negative texts of different methods are similar and not complementary. 2) It is most important to break the bias that all input texts are positive. Adding more or less negative texts should act similarly in terms of breaking the bias.
@@ 6.3.2 Subject studies on negative texts
| Whole OmniLabel | OmniLabel-Negative | D (default) | D (by length of texts) | ||||||||||||
@@ 6.3.3 Visualizations of negative texts
Fig. 14 provides good cases of negative texts from three methods that adopt LLMs. As shown, LLMs can replace objects, attributes, or relationships with other reasonable concepts. Note that negative texts from our in-context summary include the swap of concepts, where no words were removed or added. Since such swap is a feature of Winoground [45], we can conclude that the in-context summary is an effective way to learn specific patterns of a given dataset without the need of human curated promptss.
@@ 8 Implementation details
For finetuning of detectors, we initialize GLIP or FIBER models with their official checkpoints. We finetune both GLIP and FIBER a batch size of 16 and a learning rate of 1e for 1 epoch. Other hyper-parameters are kept the same as the official training. The finetuning takes around 14 hours for FIBER and 18 hours for GLIP.

### v2

@@ Abstract
The recent progress in language-based open-vocabulary object detection can be largely attributed to finding better ways of leveraging large-scale data with free-form text annotations. Training such models with a discriminative objective function has proven successful, but requires good positive and negative samples. However, the free-form nature and the open vocabulary of object descriptions make the space of negatives extremely large. Prior works randomly sample negatives or use rule-based techniques to build them. In contrast, we propose to leverage the vast knowledge built into modern generative models to automatically build negatives that are more relevant to the original data. Specifically, we use large-language-models to generate negative text descriptions, and text-to-image diffusion models to also generate corresponding negative images. Our experimental analysis confirms the relevance of the generated negative data, and its use in language-based detectors improves performance on two complex benchmarks. Code is available at https://github.com/xiaofeng94/Gen-Enhanced-Negs.
@@ 1 Introduction
Using natural language in object detection to describe semantics bears the potential to significantly increase the size of the detector’s label space and enable novel applications. While standard detectors operate on a fixed label space [39, 24, 43], natural language allows for a broad spectrum of object descriptions, ranging from generic terms like "vehicle" to specific expressions like "the red sports car parked on the left side" [42, 54, 13, 26, 55, 31, 18]. Several works advanced language-based object detection over the past few years with novel training strategies [6, 58, 4, 20, 23, 35, 33] and model architectures [16, 34, 46, 12].
@@
Our experiments demonstrate clear accuracy gains on two challenging benchmarks, +2.9AP on OmniLabel [42] and +3.3AP on D3 [54], when adding our automatically-generated negative data into the training of baseline models like GLIP or FIBER. Moreover, we provide an in-depth analysis of the generated data (text and images) and how they contribute to better language-based detection.
@@ 2 Related work
Negative samples for object detection:
The notion of negatives is crucial for training discriminative models [40, 25]. Also for object detection, hard negative mining [45] has proven beneficial for model training. However, these prior works aim to find hard negative training examples rather than negatives in the label space, because the label space is fixed in standard detection. For language-based datasets, the space of potential negatives is extremely large because object descriptions are free-form text. Prior works [5, 30, 56, 47, 44] investigate negative texts for general vision & language models with different strategies, including changing individual words (foil) with rules based on knowledge graphs [32] or with LLMs. SugarCrepe [9] shares a similar idea as us to get negative texts with in-context learning but for image-text level pretraining. For language-based detection, [8] explores such rule-based foils, while [21] uses LLMs with specific templates to replace object names with alternative descriptions. In contrast, our work (1) focuses on the localization task, (2) explores more comprehensive strategies to generate negatives with LLMs, and (3) proposes to also generate corresponding negative images with text-to-image diffusion models.
@@ 3.2.1 Generating negative descriptions with LLMs
Prior work [44, 5, 8] explored rule-based approaches to generate negative text. However, such rules are typically limited to simple knowledge graphs and are limited to replacing only individual words, often just nouns, or swapping words. In contrast, we explore more powerful LLMs to automatically generate relevant negatives. To make the negative text generation efficient and economic, in all cases, we first leverage a strong instruction-tuned LLM [36] to generate 50k positive-negative pairs, and then finetune a LLaMA-7B [48] model with those pairs to then generate negative captions on large grounding datasets. In the following, we describe three ways to instruct an LLM for generating positive-negative pairs of object descriptions:
@@
We first prompt an instruction-tuned LLM [36] to find concepts (i.e., objects, attributes and relationships) in object descriptions. Compared to rule-based parsers [52], LLMs can provide richer information. For example, for the caption “A transportation vehicle is carrying a crowd of people who are sitting and standing.”, the parser ignores “sitting” and “standing”, while LLMs regard them as attributes. Then, we pick one concept from the first step sequentially and prompt LLMs again to generate a negative caption by changing the concept. For both steps, the prompts are manually curated with the task definition and step-by-step instructions for the generation. Please find the exact prompt for the LLM in the supplement.
@@
Next, we give the LLM more freedom in generating negative descriptions. We first prompt the LLM to identify all objects in the original caption, and then to re-combine them to create a new sentence different from the original one. We allow the LLM to ignore, change or add new objects. For example, given the caption “A boy is playing with his dog” and two objects “boy” and “dog”, the LLM can output “The girl and her dog are playing fetch in the park”. Detailed prompts for both identifying objects and re-combination are in the supplement.
@@
Third, we enable LLMs to learn how to generate negative descriptions by providing human-annotated positive-negative pairs as in-context samples. We randomly sample 80 pairs of positive and negative texts from the Winoground dataset [47] and prompt the instruction-tuned LLM [36] to summarize the difference of those pairs in plain text. Then, instead of manually creating prompts to generate positive-negative pairs, we leverage the summary together with three randomly sampled Winoground pairs as prompts to the LLM, and generate several positive-negative pairs to finetune LLaMA. After finetuning, the LLaMA model is used to generate negative texts for given descriptions. This pipeline does not require hand-crafted prompts to LLMs as the explanation of the concept of negatives and how to create them. The supplement contains full prompts for generating a summary, and generating positive-negative pairs for finetuning.
@@ 3.3 Learning from negative samples
The generated data does not prescribe any specific architecture for the detector. A common choice, which we also use for our experiments, is [20, 4]. The inputs are image and text , and the output is a set of bounding boxes with corresponding logits . Here, is the number of tokens required to represent text . The ground truth can be represented by a binary assignment matrix . Rows refer to ground truth boxes and columns to tokens in . Each element indicates if a token corresponds to a box , which is given by the ground truth indices . To define a loss, bipartite graph matching associates predictions with ground truth. For matched predictions, the target vector is the corresponding row from , while it is an all-zero vector for unmatched targets. The loss is then computed as where FL refers to a focal loss. Fig. 4 illustrates the loss.
@@ 4.1 Experimental design
We choose two recently proposed benchmarks, OmniLabel [42] and D3 [54], as our test beds. These benchmarks evaluate more aspects of language-based detection than existing referring expressions [55, 31, 51] or open-vocabulary detection [18, 7] benchmarks. Specifically, both benchmarks contain complex object descriptions that go beyond simple category names from open-vocabulary detection benchmarks. Moreover, the descriptions can refer to zero, one or multiple instances in the image, in contrast to standard referring expression benchmarks. These properties enable a more stringent evaluation metric as in object detection, which is based on average precision (AP) in both OmniLabel [42] and D3 [54]. Both benchmarks provide more fine-grained metrics. OmniLabel evaluates separately for categories, descriptions, and descriptions referring to at least one object, with APc, APd and APd-P, respectively. D3 differentiates descriptions on absence (“Abs”) and presence (“Pres”) that indicate whether or not they contain any form of negation (e.g., “without”), as well as on text lengths. Finally, we create a specific split for OmniLabel, “OmniLabel-Negative”, to evaluate the model only on images that contain at least one negative description (i.e., not referring to any object).
@@
| D3 (default) | D3 (by length of texts) | ||||||
@@ 4.2 Benchmark comparisons
Tabs. 1 and 2 evaluate the impact of our generated negative training data on the OmniLabel [42] and D3 [54] benchmarks. In both tables, the first set of rows are baselines provided by the benchmarks. The following rows show the main comparisons for GLIP-T [20] and FIBER-B [4] with and without adding our generated negative training data. First, we can see that adding negative data improves results across all metrics for both models and both benchmarks. On OmniLabel, we can see a +2.9% and +2.4% increase in AP for GLIP-T and FIBER-B, respectively. Similarly, we observe a +2.3% and +3.3% increase in the main metric of D3 (AP on full descriptions) for GLIP-T and FIBER-B.
| Whole OmniLabel | OmniLabel-Negative | D3 | |||||||||
@@ 4.3 Analysis on negative texts
Effectiveness of different negative texts: We finetune FIBER-B without and with different kinds of negative texts mentioned in Sect. 3.2.1, i.e., Rule-based foils, LLM-based foils, Re-combination with LLMs, In-context summary with LLMs, and present results in Table 4. We find all kinds of negatives improve the original FIBER-B on both OmniLabel and D3 benchmarks. Negative texts from LLMs generally achieve better results compared to LLM-based foils, which indicates that LLMs are powerful tools for negative text generation. Moreover, both recombination and in-context summary with LLMs outperform LLM-based foils in all metrics except APd-P. Note that APd-P refers to evaluations without negative label spaces, which is a task weaker than language-based detection. Based on such results, we argue that although word foils provide promising results in traditional studies [8, 44], it is sub-optimal to LLMs. We need to explore varied ways to unlock the ability of LLMs. We believe that our two solutions, i.e., Re-combination and In-context summary, provide a good starting point for future studies. Besides using only one kind of negative texts, we also explore the combinations of different kinds of negative texts in the supplement.
@@ 4.4 Analysis on negative images
Following the idea of concatenating the positive and negative captions as the text input, we concatenate the positive and negative images as one input image during training. See supplement for an example. In this way, models are forced to tell the difference between the positive and negative images within one training iteration, which helps detectors to learn better about the negative. As shown in Table 4, such a simple technique improves upon “+ Box&CLIP filters” both on OmniLabel and D3. Furthermore, we ensemble the weights of two FIBER models, one finetuned with negative texts only, and the other finetuned with both negative texts and images. Finally, compared to using negative texts only, we gain 1.3 APd on OmniLabel and no performance drop on D3.
@@
Table 4 shows that negative images help on OmniLabel but not much on D3. We explore this on a data basis. We first crop image regions for generated images, OmniLabel images, D3 images, and Flickr30k images based on the bounding boxes. Then, we randomly select 1000 image regions and feed them into a CLIP image encoder to get CLIP embeddings. Later, we input those embeddings to t-SNE [49] to illustrate the similarities between different image regions. As shown in Fig. 11, D3’s regions are grouped into several clusters, while OmniLabel and our generated regions are scattered in the center. This indicates that there is a clear domain gap between D3 and the others. Thus, it is plausible that our generated images only helps on OmniLabel. In our view, the gap comes from that D3 collect data in groups based on categories. In contrast, OmniLabel collects data randomly.
@@ 5 Conclusion
Acknowledgments: This research project has been partially funded by research grants to Dimitris N. Metaxas through NSF: 2310966, 2235405, 2212301, 2003874, and FA9550-23-1-0417.
@@ 6.2 Prompts to LLMs for negative text generation
As mentioned in the main paper, we first leverage a strong instruction-tuned LLM [36] to generate 50k positive-negative pairs, and then finetune a LLaMA-7B [48] model with those pairs to later generate negative captions on large grounding datasets. We describe prompts we used to generate positive-negative pairs for finetuning as follows.
@@ 6.3.1 Combining negative texts of different sources
In this part, we explore to combine negative texts of different sources. When two or more sources are combined, we randomly sample negative texts from the combined source during training. Table 5 provides the results of different combinations, from which we have the following findings. First, all combinations outperform the original FIBER in terms of almost all the metrics, which indicates the effectiveness of adding negative texts. Second, there is no combination that consistently outperforms others in all metrics. For example, “+ (4) Texts for negative images” achieves the leading performance on the whole OmniLabel but not on OmniLabel-Negative and D3. Moreover, the performance gaps between different combinations are minor. Some possible explanations to such results are: 1) All four methods generate negative texts mainly by changing concepts of the original captions. Thus, negative texts of different methods are similar and not complementary. 2) It is most important to break the bias that all input texts are positive. Adding more or less negative texts should act similarly in terms of breaking the bias.
@@ 6.3.2 Subject studies on negative texts
| Whole OmniLabel | OmniLabel-Negative | D3 (default) | D3 (by length of texts) | ||||||||||||
@@ 6.3.3 Visualizations of negative texts
Fig. 14 provides good cases of negative texts from three methods that adopt LLMs. As shown, LLMs can replace objects, attributes, or relationships with other reasonable concepts. Note that negative texts from our in-context summary include the swap of concepts, where no words were removed or added. Since such swap is a feature of Winoground [47], we can conclude that the in-context summary is an effective way to learn specific patterns of a given dataset without the need of human curated prompts.
@@ 8 Implementation details
For finetuning of detectors, we initialize GLIP or FIBER models with their official checkpoints. We finetune both GLIP and FIBER a batch size of 16 and a learning rate of 1e-5 for 1 epoch. Other hyper-parameters are kept the same as the official training. The finetuning takes around 14 hours for FIBER and 18 hours for GLIP.